# Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method

### v4

@@ Abstract
Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method
Zitian Gao1, Yihao Xiao2 1School of Computer Science, The University of Sydney, Australia 2School of Finance, Shanghai University of Finance and Economics, China zgao3186@uni.sydney.edu.au, 2023121716@stu.sufe.edu.cn
@@ 1 Introduction
Previous work has primarily focused on simple deep learning and machine learning methods, which often rely on limited financial data or single time series analyses Wang and Ihlamur (2024); Lewis et al. (2021). Specifically, most models focus on using historical financial indicators, market trends, or other single-dimensional data to predict the future performance of startups. While these models can capture the growth potential of startups in some cases, they often overlook critical factors within the broader ecosystem in which startups operate, especially the competitive and cooperative relationships between companies. Startups typically operate in environments with limited resources, and their interactions with one another play a significant role in their success or failure. Traditional methods have failed to fully leverage this relational data, leading to poor performance in predicting startup success, particularly in cases with sparse data and low signal strength. These models struggle to accurately identify early success signals under such conditions.
@@
Startups have various potential development goals, with the most desirable being successful acquisition by a large company or going public. While many studies and VCs focus on predicting whether a company will successfully go public as a binary classification task Cao et al. (2022); Wang and Ihlamur (2024), we attempt to predict what will happen to the company next. This approach aims to provide VCs with opportunities not only to consider high-risk, high-reward companies but also to build lower-risk investment portfolios and determine when to exit Bienz and Walz (2010); Schwienbacher (2008).
@@
In contrast to another common direction in financial time series prediction—stock trend forecasting—fine-grained, structured time series data used for predicting stock trends is readily available Sawhney et al. (2021); Wang et al. (2021, 2022). By comparison, predicting the success of startups in the venture capital (VC) field requires forecasting the success of unlisted startups with minimal available data, often relying on hard-to-access unstructured text such as team backgrounds, patents, and product reviews.
@@ 2.1 Startup Success Prediction
Predicting the success of startups has garnered increasing attention from scholars, especially in the venture capital field. Existing research primarily focuses on developing methods to represent company features and relationships and on building predictive models using machine learning and deep learning techniques. For example, Cao et al. (2024) constructed CompanyKG to represent and learn various features and relationships between companies, aiming to capture the impact of these features on a company’s success. Żbikowski and Antosiuk (2021) used Crunchbase data to create an XGBoost-based model for predicting the likelihood of a company’s success. Kim et al. (2020) used a doc2vec deep learning method to extract feature vectors from startup profiles and patent abstracts, applying them to filter promising startups in certain fields.In addition, Ross et al. (2021) introduced the CapitalVX framework, which leverages general business information of startups to predict optimal entry and exit timings, providing valuable insights into market dynamics and investment strategies. Although these approaches have made some progress, they still face challenges in handling complex nonlinear relationships and long-term predictions. Lyu et al. (2021) designed an incremental representation learning mechanism and a sequential learning model based on graph neural networks to predict the success of startups. Wang et al. (2020) performed preprocessing and exploratory analysis on Kickstarter data and then introduced deep learning algorithms like multilayer perceptrons for predicting crowdfunding outcomes. To address long-term revenue prediction, Cao et al. (2022) proposed the Simulation of Information Revenue Extrapolation (SiRE) algorithm to generate fine-grained long-term revenue forecasts.Furthermore, some studies, such as Leogrande et al. (2021), have explored the use of unstructured data, including news articles and social media content, for predicting startup success, highlighting the potential of diverse data sources in this domain.
@@ 2.2 Retrieval-Augmented Generation In Finance
In the finance domain, Retrieval-Augmented Generation Lewis et al. (2021) has become a significant method, particularly for answering questions based on large datasets. RAG methods extract information from retrievable text regions, providing a solid foundation for generation tasks. For instance, Wang and Ihlamur (2024) developed a startup success prediction framework (SSFF) that incorporates RAG as an external knowledge module to gain a comprehensive understanding of market landscapes. However, in more complex scenarios, such as handling multi-dimensional data and intricate relationships, the capabilities of traditional RAG may be limited.
@@ 3 Methodology
Our approach consists of two stages: First, we input a large amount of unstructured news text into GraphRAG, where during the clustering and retrieval process, global statistics are dynamically updated to accurately cluster relationships between entities. This results in a directed knowledge graph of companies, with edges representing both direct and indirect relationships between companies. In the second stage, inspired by Ibrahim et al. (2022) and Barigozzi and Brownlees (2019), we transform the knowledge graph into a mask matrix using the Leiden algorithm Traag et al. (2019). This matrix is then used as a regularizer in the multivariate time series model, combined with relevant covariates. This method effectively enhances the generalization ability of models on scarce company data and significantly improves prediction performance, as illustrated in Figure 1. The specific algorithm details are provided in Algorithms 1 and 2.
@@ 3.1 Sequence-to-Sequence Representation
The output sequence $Y_{n}$ represents the company’s post-IPO performance over several quarters, specifically in terms of the price-to-book ratio (P/B ratio). The P/B ratio, defined as the ratio of the company’s stock price (Price) to its book value (Book Value), serves as a crucial metric in venture capital. Unlike the price-to-earnings ratio (P/E ratio), which can be volatile in the short term due to fluctuations in earnings, the P/B ratio offers a more stable measure of a company’s intrinsic value and long-term sustainability Nissim and Penman (2003). By predicting the P/B ratio over multiple quarters, our model provides a detailed projection of a company’s ongoing financial health and shareholder profitability.

### v5

# soft open fences

@@ 1 Introduction
Previous work has primarily focused on simple deep learning and machine learning methods, which often rely on limited financial data or single time series analyses Wang & Ihlamur (2024); Lewis et al. (2021). Specifically, most models focus on using historical financial indicators, market trends, or other single-dimensional data to predict the future performance of startups. While these models can capture the growth potential of startups in some cases, they often overlook critical factors within the broader ecosystem in which startups operate, especially the competitive and cooperative relationships between companies. Startups typically operate in environments with limited resources, and their interactions with one another play a significant role in their success or failure. Traditional methods have failed to fully leverage this relational data, leading to poor performance in predicting startup success, particularly in cases with sparse data and low signal strength. These models struggle to accurately identify early success signals under such conditions.
@@
Startups have various potential development goals, with the most desirable being successful acquisition by a large company or going public. While many studies and VCs focus on predicting whether a company will successfully go public as a binary classification task Cao et al. (2022); Wang & Ihlamur (2024), we attempt to predict what will happen to the company next. This approach aims to provide VCs with opportunities not only to consider high-risk, high-reward companies but also to build lower-risk investment portfolios and determine when to exit Bienz & Walz (2010); Schwienbacher (2008).
@@
In contrast to another common direction in financial time series prediction—stock trend forecasting—fine-grained, structured time series data used for predicting stock trends is readily available Sawhney et al. (2021); Wang et al. (2021; 2022). By comparison, predicting the success of startups in the venture capital (VC) field requires forecasting the success of unlisted startups with minimal available data, often relying on hard-to-access unstructured text such as team backgrounds, patents, and product reviews.
@@ 2.1 Startup Success Prediction
Predicting the success of startups has garnered increasing attention from scholars, especially in the venture capital field. Existing research primarily focuses on developing methods to represent company features and relationships and on building predictive models using machine learning and deep learning techniques. For example, Cao et al. (2024) constructed CompanyKG to represent and learn various features and relationships between companies, aiming to capture the impact of these features on a company’s success. Żbikowski & Antosiuk (2021) used Crunchbase data to create an XGBoost-based model for predicting the likelihood of a company’s success. Kim et al. (2020) used a doc2vec deep learning method to extract feature vectors from startup profiles and patent abstracts, applying them to filter promising startups in certain fields.In addition, Ross et al. (2021) introduced the CapitalVX framework, which leverages general business information of startups to predict optimal entry and exit timings, providing valuable insights into market dynamics and investment strategies. Although these approaches have made some progress, they still face challenges in handling complex nonlinear relationships and long-term predictions. Lyu et al. (2021) designed an incremental representation learning mechanism and a sequential learning model based on graph neural networks to predict the success of startups. Wang et al. (2020) performed preprocessing and exploratory analysis on Kickstarter data and then introduced deep learning algorithms like multilayer perceptrons for predicting crowdfunding outcomes. To address long-term revenue prediction, Cao et al. (2022) proposed the Simulation of Information Revenue Extrapolation (SiRE) algorithm to generate fine-grained long-term revenue forecasts.Furthermore, some studies, such as Leogrande et al. (2021), have explored the use of unstructured data, including news articles and social media content, for predicting startup success, highlighting the potential of diverse data sources in this domain.
@@ 2.2 Retrieval-Augmented Generation In Finance
In the finance domain, Retrieval-Augmented Generation Lewis et al. (2021) has become a significant method, particularly for answering questions based on large datasets. RAG methods extract information from retrievable text regions, providing a solid foundation for generation tasks. For instance, Wang & Ihlamur (2024) developed a startup success prediction framework (SSFF) that incorporates RAG as an external knowledge module to gain a comprehensive understanding of market landscapes. However, in more complex scenarios, such as handling multi-dimensional data and intricate relationships, the capabilities of traditional RAG may be limited.
@@ 3 Methodology
Our approach consists of two stages: First, we input a large amount of unstructured news text into GraphRAG, where during the clustering and retrieval process, global statistics are dynamically updated to accurately cluster relationships between entities. This results in a directed knowledge graph of companies, with edges representing both direct and indirect relationships between companies. In the second stage, inspired by Ibrahim et al. (2022) and Barigozzi & Brownlees (2019), we transform the knowledge graph into a mask matrix using the Leiden algorithm Traag et al. (2019). This matrix is then used as a regularizer in the multivariate time series model, combined with relevant covariates. This method effectively enhances the generalization ability of models on scarce company data and significantly improves prediction performance, as illustrated in Figure 1. The specific algorithm details are provided in Algorithms 1 and 2.
@@ 3.1 Sequence-to-Sequence Representation
The output sequence $Y_{n}$ represents the company’s post-IPO performance over several quarters, specifically in terms of the price-to-book ratio (P/B ratio). The P/B ratio, defined as the ratio of the company’s stock price (Price) to its book value (Book Value), serves as a crucial metric in venture capital. Unlike the price-to-earnings ratio (P/E ratio), which can be volatile in the short term due to fluctuations in earnings, the P/B ratio offers a more stable measure of a company’s intrinsic value and long-term sustainability Nissim & Penman (2003). By predicting the P/B ratio over multiple quarters, our model provides a detailed projection of a company’s ongoing financial health and shareholder profitability.
@@ 5 Limitation
6 Future Work
While our proposed GraphRAG-enhanced multivariate time series analysis method has shown promise in predicting startup success, there are several areas that warrant further investigation to address current limitations and enhance the model’s capabilities.
Generalization Across Datasets: A key limitation of our current work is the reliance on non-structured data, which can be difficult to obtain consistently across different regions and industries. The dataset used in this study is primarily focused on the Chinese A-share market, which may limit the generalizability of our findings. Future research should focus on testing the model’s performance on datasets from other countries and markets to evaluate its robustness and adaptability across different economic and cultural environments.
Exploration of Soft Masking: The current approach utilizes a specific method for constructing the mask matrix based on the Leiden algorithm. However, there is potential to explore other soft masking techniques that might offer improved performance. Future work could investigate the use of different clustering algorithms to generate mask matrices, potentially enhancing the model’s ability to capture complex inter-company relationships.